Find the correct solution set that satisfies the inequality
5x+2≥7(x+2)

Answer by Theo(13342)

5x + 2 >= 7 * (x + 2)

simplify to get 5x + 2 >= 7x + 14

subtract 5x from both sides of the equation and subtract 14 from both sides of the equation to get 2 - 14 >= 7x - 5x.

combine like terms to get -12 >= 2x

divide both sides of the equation by 2 to get -6 >= x

this is the same as x <= -6.

that's your solution.

you could also have solved it this way:

start with 5x + 2 >= 7x + 14

subtract 7x from both sides of the equation and subtract 2 from both sides of the equation to get 5x - 7x >= 14 - 2

combine like terms to get -2x >= 12

divide both sides of the equation by -2 to get x <= -6.

note that multiplying or dividing both sides of an inequality reverses the inequality.

example:

5 > 3
multiply both sides by -1 gets you -5 < -3

since division is just another form of multiplication, it works for division as well.

example:

10 < 20
divide both sides by -2 gets you -5 > -10

division by -2 is the same as multiplication by -1/2

10 < 20
multiply both sides by -1/2 gets you -5 > -10.

anyway, back to the problem at hand.

your solution is the x <= -6.

to confirm replace x in the original equation with -8, -6, -4.

original equation is 5x + 2 >= 7 * (x + 2)

when x = -8, this becomes -38 >= -42 which is true.
when x = -6, this becomes -28 >= -28 which is true.
when x = -4, this becomes -18 >= -14 which is false.

solution is confirmed to be good because, when x <=-6, the original equation is true, and when x > -6, the original equation is false.